OpenStudy (anonymous):

|-11| + 52 = 50y - |-31| - 49y

OpenStudy (anonymous):

y=94

OpenStudy (anonymous):

look guys

OpenStudy (accessdenied):

Oh! You should have written "5^2" to be clear on that squared term. :)

OpenStudy (accessdenied):

|-11| = 11 5^2 = 25 |-31| = 31 11 + 25 = 50y - 31 - 49y 36 = y - 31 67 = y

OpenStudy (anonymous):

lol your right